OpenStudy (anonymous):

find an equation of the line containing the given pair of points (1,5) & (3,6). y=

OpenStudy (anonymous):

y=ax+b \[a = \Delta Y / \Delta X\] = (6-5)/(3-1) = 1/2=0.5 so y=0.5x + b since y=5 and x=1: 5 = 0.5 * 1 + b which gives b=4.5 So y=0.5x+4.5 Check: 6=0.5 * 3 +b which gives b=4.5 too, so it's correct!